# Pragmatic Curiosity: A Hybrid Learning-Optimization Paradigm via Active Inference

### v1

@@ 3.1 Active Inference as Expected Free Energy Minimization
$\begin{split}G&=-\mathbb{E}_{q(y|x)}\left[D_{\text{KL}}(q(s|x,y)||q(s))\right]-\mathbb{E}_{q(y|x)}\log p(y)\\ &=\underbrace{-I_{q}(s;(x,y))}_{\text{epistemic}}\underbrace{-\mathbb{E}_{q(y|x)}\log p(y)}_{\text{pragmatic}},\end{split}$ (3)
where $I_{q}(\cdot)$ represents the mutual information given the surrogate model $q(\cdot)$.
Proof.
See Appendix B. ∎
We can see that by construction, EFE strikes a balance between information-seeking and goal-directed behavior under some prior preferences. It bounds the difference between epistemic value (about parameters) and pragmatic value (about outcomes), which captures the imperative to maximize the epistemic value (i.e., information gain about latent states), from interactions with the environment, while maximizing the pragmatic value (i.e., expected preference alignment), regarding prior preferences. This crucial aspect of AIF effectively addresses the “explore-exploit dilemma” because the imperatives for exploration and exploitation are just two aspects of EFE:
Pragmatic Value (Exploitation): This term encourages goal-directed behavior by favoring actions expected to yield preferred outcomes. Encoded by a prior distribution over desired observations, it functions similarly to a utility or reward function in reinforcement learning (Millidge et al., 2020), driving the agent to exploit its current knowledge to achieve its goals.
Epistemic Value (Exploration): This term promotes information-seeking behavior by favoring actions expected to reduce uncertainty about the underlying system maximally. It quantifies the expected information gain about the model’s parameters, driving the agent to explore the environment to refine its world model.
3.2 Reinterpretation of Acquisition Strategies in BO and BED
Crucially, minimizing EFE serves as a unifying umbrella principle. Many classic acquisition strategies in BO and BED can be reinterpreted as special cases of minimizing EFE, as shown in Table 1.
@@ 3.2 Reinterpretation of Acquisition Strategies in BO and BED
Acquisition Strategy Acquisition Function $\alpha(x)$ Expected Free Energy
Parameters $s$ Preferences $p(y)$
Expected Information Gain (Chaloner and Verdinelli, 1995) $I(\theta;(x,y)|\mathcal{D}_{t})$ $\theta$ -
Entropy Search (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014) $I(x^{*};(x,y)|\mathcal{D}_{t})$ $x^{*}$ -
Max-value Entropy Search (Wang and Jegelka, 2017) $I(y^{*};(x,y)|\mathcal{D}_{t})$ $y^{*}$ -
Joint Entropy Search (Hvarfner et al., 2022b) $I((x^{*},y^{*});(x,y)|\mathcal{D}_{t})$ $(x^{*},y^{*})$ -
Bayesian Algorithm Execution (Neiswanger et al., 2021) $I(\mathcal{O}_{\mathcal{A}}(f);(x,y)|\mathcal{D}_{t})$ $\mathcal{O}_{\mathcal{A}}(f)$ -
GP-Upper Confidence Bound $\mu_{t}(x)+\beta^{1/2}\sigma_{t}(x)$ $f_{\mathcal{X}}$ $\exp\{y\}$
@@
The reinterpretations of most of the acquisition strategies in Table 1 are straightforward according to their definitions. However, placing a rather intuitive GP-UCB strategy within this framework seems implicit. To reveal their connection, we rely on the following lemma:
Lemma 3.2.
Let $\mathbf{X}\subseteq\mathcal{X}$ be a subset of inputs, and the corresponding function values evaluated at those inputs be denoted as $f_{\mathbf{X}}$. Given a historical dataset $\mathcal{D}_{t}$, and new measurements $\mathbf{Y}$ observed at $\mathbf{X}$, the mutual information
$I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})|\mathcal{D}_{t})=I(f_{\mathbf{X}};\mathbf{Y}|\mathcal{D}_{t}),$
@@
See Appendix C. ∎
Then if we assume constant Gaussian noises $\mathcal{N}\sim(0,\sigma^{2})$ for the observations, we have
$\begin{split}I(f_{x};y|\mathcal{D}_{t})&=H(y|\mathcal{D}_{t})-H(y|f_{x},\mathcal{D}_{t})\\ &=\frac{1}{2}\log(1+\sigma^{-2}\sigma_{t}^{2}(x)),\end{split}$
If we choose $\beta=\frac{1}{2}\log(1+\sigma^{-2})$, then the epistemic term in EFE, i.e., $I(f_{x};y|\mathcal{D}_{t})$, provides an upper bound of the square of the exploration term $\beta^{1/2}\sigma_{t}(x)$ in GP-UCB.
This reveals the close relationship between GP-UCB and AIF, showing that even a seemingly pure intuitive strategy, like GP-UCB, can have rather rigorous mathematical foundations underlying it.
4 A New Paradigm to Derive Acquisition Functions
The preference distribution $p(y)$ in Table 1 can be interpreted as a softmax transformation of a time-varying value function over the outcomes. Generalizing from this observation, we propose a new paradigm for deriving acquisition functions applicable to a wider class of hybrid problems other than traditional BO and BED.
Grounded on Boltzmann distribution (also called Gibbs distribution) in statistical mechanics, we define an Boltzmann operator $\mathcal{B}_{\beta}$ that maps an energy function $h(z)$ over domain $\mathcal{Z}$ into a Boltzmann distribution:
$(\mathcal{B}_{\beta}[h])(z)\coloneq\frac{e^{-h(z)/\beta}}{\int_{\mathcal{Z}}e^{-h(z)/\beta}dz},$
where $\beta$ is called a temperature parameter (in allusion to statistical mechanics). A higher temperature results in a more uniform output distribution (i.e.“more random” with higher entropy), while a lower temperature results in a sharper output distribution, with the maximizers of $h(z)$ dominating.
Then, for any pre-defined energy function $h(y|\mathcal{D}_{t})$, we can transfer it into a Boltzmann distribution and use it as the preference distribution $p(y)$ in (3):
$\begin{split}G&=-I(s;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[\log{e^{-h(y|\mathcal{D}_{t})/\beta}}]+Z\\ &=-I(s;(x,y)|\mathcal{D}_{t})+\frac{1}{\beta}\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[h(y|\mathcal{D}_{t})]+Z,\end{split}$
where $p(y|x,\mathcal{D}_{t})$ is the predictive distribution of a surrogate model constructed from the historical data $\mathcal{D}_{t}$, and $Z=\log\int_{\mathcal{Y}}e^{-h(y|\mathcal{D}_{t})/\beta}dy$ is a normalization constant independent of $y$.
Therefore, we propose a new acquisition paradigm, pragmatic curiosity, by minimizing the EFE, i.e.,
$x_{t+1}=\arg\max_{x\in\mathcal{X}}\{\beta_{t}I(s;(x,y)|\mathcal{D}_{t})-\mathbb{E}_{p(y|x,\mathcal{D}_{t})}[h(y|\mathcal{D}_{t})]\},$
where $\beta_{t}\geq 0$ is a parameter that modulates the degree of curiosity, and $h(y|\mathcal{D}_{t})$ is a problem-dependent energy function that captures the goal (in terms of regret w.r.t. a certain outcome $y$).
The curiosity $\beta_{t}$ plays an important role in balancing the performance between learning and optimization. Li et al. (2026) derive a formal theoretical guarantee of posterior consistency and bounded cumulative regret under a sufficient curiosity condition, and provide practical design guidelines for choosing $\beta_{t}$ in different settings.
This paper instead examines how a problem-dependent energy function, $h(y|\mathcal{D}_{t})$, provides a flexible mechanism for explicitly representing uncertainty over the goals themselves. This added expressiveness enables acquisition design for a broad class of complex settings that standard methods often overlook, including tasks with evolving goals (where conditions change over time) and implicit goals (not specified a priori). The effectiveness and versatility of this formulation on such tasks are demonstrated in the next section.

### v2

@@ 3.2 A Principled Decision Paradigm for Hybrid Learning and Optimization
$\begin{split}\text{Term 1}&=-\mathbb{E}_{q(y,s\mid x)}\left[\log p(s\mid x,y)-\log q(s\mid x,y)+\log q(s\mid x,y)-\log q(s\mid x)\right]\\ &=-\mathbb{E}_{q(y\mid x)}\mathbb{E}_{q(s\mid x,y)}\left[\log p(s\mid x,y)-\log q(s\mid x,y)+\log q(s\mid x,y)-\log q(s\mid x)\right]\\ &=-\mathbb{E}_{q(y\mid x)}\left[\mathbb{E}_{q(s\mid x,y)}\left[\log p(s\mid x,y)-\log q(s\mid x,y)\right]+\mathbb{E}_{q(s\mid x,y)}\left[\log q(s\mid x,y)-\log q(s\mid x)\right]\right]\\ &=-\mathbb{E}_{q(y\mid x)}\left[\underbrace{-D_{\text{KL}}(q(s\mid x,y)\|p(s\mid x,y))}_{\text{Term 3}}+\underbrace{D_{\text{KL}}(q(s\mid x,y)\|q(s\mid x))}_{\text{Term 4}}\right].\end{split}$
Since the surrogate belief over $s$ is not updated until an observation is obtained, it does not depend on $x$ alone; hence $q(s\mid x)=q(s)$. Therefore, Term 4 becomes $D_{\text{KL}}(q(s\mid x,y)\|q(s))$, which is precisely the epistemic value: the information gained about the latent quantity $s$ from a prospective observation $(x,y)$.
By contrast, Term 3 remains intractable because it depends on the true posterior $p(s\mid x,y)$. The best the agent can access is its current surrogate belief constructed from the historical data $\mathcal{D}_{t}$, namely $q(\cdot)=p(\cdot\mid\mathcal{D}_{t})$. This motivates the first approximation: we treat the surrogate belief as the best currently available approximation to the true model, i.e., $q(s)\approx p(s)$. If both the surrogate and true models update according to Bayes’ rule, $q(s\mid x,y)=\frac{p(x,y\mid s)q(s)}{\int p(x,y\mid s)q(s)ds},\quad p(s\mid x,y)=\frac{p(x,y\mid s)p(s)}{\int p(x,y\mid s)p(s)ds},$ then Term 3 vanishes under this approximation, and the resulting decision rule is Bayes-optimal with respect to the agent’s current belief. This is the only legitimate basis for action: a decision-maker cannot optimize with respect to an inaccessible ground truth, but only with respect to its present posterior approximation. Moreover, under standard consistency conditions, the discrepancy between surrogate and truth shrinks as more data are collected.
The second obstacle lies in Term 2, which depends on the true predictive distribution $p(y\mid x)$. We therefore introduce a second approximation: we replace $p(y\mid x)$ by a preference distribution $p_{\text{pref}}(y)$ that does not depend on $x$. This converts Term 2 into $\mathbb{E}_{q(y\mid x)}\log p_{\text{pref}}(y)$, which evaluates preferred outcomes under the predictive model induced by choosing $x$. Intuitively, the agent now asks: if I take action $x$, what outcomes am I likely to see, and how desirable are those outcomes?
To define a preference distribution from a task-level notion of desirability, we introduce a Boltzmann operator $\mathcal{B}_{\beta}$ that maps a nonnegative potential energy function $h:\mathcal{Z}\to\mathbb{R}_{\geq 0}$ into a probability distribution:
$(\mathcal{B}_{\beta}[h])(z)\coloneq\frac{e^{-h(z)/\beta}}{\int_{\mathcal{Z}}e^{-h(z)/\beta}dz}.$
This construction is inspired by the Boltzmann (or Gibbs) distribution in statistical mechanics. The parameter $\beta$ plays the role of temperature: larger $\beta$ yields a flatter, higher-entropy distribution, whereas smaller $\beta$ concentrates more sharply around the minimizers of $h$.
Applying this operator to a possibly time-varying potential energy function $h(y\mid\mathcal{D}_{t})$ gives a preference distribution over outcomes, and hence
$\begin{split}G&\approx-I(s;(x,y)\mid\mathcal{D}_{t})-\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}[\log{e^{-h(y\mid\mathcal{D}_{t})/\beta}}]+Z\\ &=-I(s;(x,y)\mid\mathcal{D}_{t})+\frac{1}{\beta}\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}[h(y\mid\mathcal{D}_{t})]+Z,\end{split}$
where $q(y\mid x,\mathcal{D}_{t})$ is the predictive distribution of a surrogate model constructed from the historical data $\mathcal{D}_{t}$, and $Z=\log\int_{\mathcal{Y}}e^{-h(y\mid\mathcal{D}_{t})/\beta}dy$ is a normalization constant independent of $x$.
This leads to our proposed decision principle:
$x_{t+1}=\arg\max_{x\in\mathcal{X}}\{\beta_{t}\underbrace{I(s;(x,y)\mid\mathcal{D}_{t})}_{\text{epistemic}}-\underbrace{\mathbb{E}_{q(y\mid x,\mathcal{D}_{t})}[h(y\mid\mathcal{D}_{t})]\}}_{\text{pragmatic}}.$
Here, the conditional mutual information $I(s;(x,y)\mid\mathcal{D}_{t})$ captures the epistemic drive to reduce uncertainty about latent quantities of interest; the expected potential $\mathbb{E}[h(y\mid\mathcal{D}_{t})]$ captures the pragmatic drive to avoid undesirable outcomes; and $\beta_{t}\geq 0$ controls the trade-off between the two.
By construction, this decision rule balances information-seeking and goal-directed behavior within a single objective. Rather than treating exploration and exploitation as competing heuristics, it expresses them as two inseparable aspects of one decision criterion: the agent seeks information insofar to improve outcome-sensitive action. In other words, it demonstrates a pragmatic curiosity (PraC).
3.3 A Unified View of Acquisition Strategies in BO and BED
PraC serves as a unifying umbrella principle for both learning-oriented and optimization-oriented acquisition strategies. By varying the epistemic target $s$ and the pragmatic potential $h_{t}(y)$, many classical acquisition rules from BO and BED can be recovered as special cases or close approximations, as summarized in Table 1.
@@ 3.3 A Unified View of Acquisition Strategies in BO and BED
Acquisition Strategy Acquisition Function Pragmatic Curiosity
$s$ $h_{t}(y)$
Expected Information Gain (Chaloner and Verdinelli, 1995b) $I(\theta;(x,y)\mid\mathcal{D}_{t})$ $\theta$ -
Entropy Search (Hennig and Christian J. Schuler, 2012; Hernández-Lobato et al., 2014) $I(x^{*};(x,y)\mid\mathcal{D}_{t})$ $x^{*}$ -
Max-value Entropy Search (Wang and Jegelka, 2017) $I(y^{*};(x,y)\mid\mathcal{D}_{t})$ $y^{*}$ -
Joint Entropy Search (Hvarfner et al., 2022) $I((x^{*},y^{*});(x,y)\mid\mathcal{D}_{t})$ $(x^{*},y^{*})$ -
Bayesian Algorithm Execution (Neiswanger et al., 2021) $I(\mathcal{O}_{\mathcal{A}}(f);(x,y)\mid\mathcal{D}_{t})$ $\mathcal{O}_{\mathcal{A}}(f)$ -
GP-Upper Confidence Bound  [^1] Strictly speaking, the correspondence for GP-UCB is approximate rather than exact, because GP-UCB uses a first-order standard-deviation term rather than a second-order variance term. Detailed derivations are provided in Appendix B. Still, the connection reveals the information-theoretic structure underlying this heuristic form. (Srinivas et al., 2009) $\mu_{t}(x)+\beta^{1/2}\sigma_{t}(x)$ $f_{\mathcal{X}}$ y-y
@@
More importantly, this unifying principle opens the door to acquisition design for a broader class of hybrid learning-and-optimization problems that go beyond classical BO and BED. Such problems are typically governed by three coupled meta-decisions: (i) what the agent seeks to know, (ii) what it seeks to achieve, and (iii) how it trades off those two. In the next section, we formalize these three design axes through symbols, regrets, and curiosity.
4 The Triad of Meta-Decisions: Symbols, Regrets, and Curiosity
One central contribution of PraC is that it foregrounds three meta-decisions underlying any hybrid learning-and-optimization problem. First, the agent must decide which latent quantity should be represented and clarified; this is implied by the choice of $s$. Second, it must decide how outcomes should be evaluated; this is encode through the choice of $h_{t}$. Third, it must decide how strongly unresolved uncertainty should influence action; this corresponds to the choice of $\beta_{t}$.
These three choices are the critical operational levers through which an abstract decision paradigm becomes a practical decision rule. This section develops those three choices both conceptually and operationally. We formalize symbols to specify the latent quantities whose clarification can change what the agent ought to do, regrets to provide an operational loss landscape over outcomes, and curiosity to regulate the exchange rate between clarifying knowledge and pursuing goals.
4.1 Symbols and Models
“All models are wrong, but some are useful.” A model becomes useful when it represents the distinctions whose clarification can change what we should do. In PraC, we call these distinctions symbols.
Different modeling regimes induce different forms of symbols.
Parametric models.
In parametric models, symbols are typically fixed global symbols: the latent state is represented by a finite-dimensional parameter or state variable whose components retain persistent semantic identities across decisions. Examples include unknown physical coefficients, latent modes, transition parameters, or class labels. In this regime, the epistemic objective is to reduce uncertainty about these stable hidden meanings.
Non-parametric models.
In non-parametric models, however, there may be no finite set of globally privileged coordinates. Instead, the relevant symbols are often induced local symbols, created by the contemplated observations themselves. This is formalized by the following representation-compression property.
Lemma 4.1 (Representation Compression).
Let $\mathbf{X}\subseteq\mathcal{X}$ be a subset of inputs, and let $f_{\mathbf{X}}$ denote the corresponding function values. Given a historical dataset $\mathcal{D}_{t}$ and new measurements $\mathbf{Y}$ observed at $\mathbf{X}$, then
$I(f_{\mathcal{X}};(\mathbf{X},\mathbf{Y})\mid\mathcal{D}_{t})=I(f_{\mathbf{X}};\mathbf{Y}\mid\mathcal{D}_{t}),$
@@ 4.1 Symbols and Models
See Appendix A. ∎
Lemma 4.1 shows that, under a non-parametric model such as a Gaussian process, the information gained from observing $\mathbf{Y}$ at $\mathbf{X}$ is mediated entirely through the function values $f_{\mathbf{X}}$ that directly generate those observations. These local evaluations play a role analogous to parameters in parametric models or latent states in partially observed dynamical systems, but they need not belong to a fixed global representation.
This compression reveals both the power and the cost of non-parametric representation. The power is parsimony: the agent need not reason about every component of an infinite-dimensional latent object when only a local slice can influence the prospective observation. The cost is semantic instability: because the symbol is query-dependent, its meaning is induced by the decision context rather than fixed in advance.
This also suggests a broader role for PraC. PraC asks not merely how expressive a model is, but whether the symbol used by the acquisition is sufficient for action evaluation, minimal enough to avoid epistemic redundancy, and structured enough to support interpretable decision-making. Thus beyond selecting actions, PraC can also guide the construction of representations. When the relevant structure is unclear, induced local symbols can reveal which distinctions are repeatedly useful for decision. Over time, such recurring local distinctions may be consolidated into more stable global symbols. In this sense, PraC does not only act within a representation; it also provides a criterion for which representations deserve to persist.
4.2 Regrets and Goals
“All successes are successful alike, all failures are failed in their own way.” Most of the time, we don’t know what the goal is until we arrive there. It seems easier to address what it is not than what it is: success is often rare and difficult to characterize in full, whereas shortfall, inconsistency, or violation is much more common and easier to recognize and certify.
For this reason, it is often more natural to encode what an outcome fails to achieve than to assume that a complete reward landscape is already known. Under this view, we define regret as the shape cast by a goal onto outcome space. Regret thus acts as a carrier of goal: it gives an operational representation of what matters by quantifying how an outcome falls short of what is desired.
We begin by formally defining a generalized notion of regret.
Definition 4.2 (Regret Function).
Let $\mathcal{Y}$ denote the outcome space, and let $\xi$ denote a desired reference, such as an optimal outcome, a target set, a safety specification, or a preference criterion. A regret function is a mapping
$r(\cdot;\xi):\mathcal{Y}\to\mathbb{R}_{\geq 0}$
that quantifies how an outcome $y$ falls short of the desired reference $\xi$. It obtains 0 for the desired outcomes, i.e., $\inf_{y\in\mathcal{Y}}r(y;\xi)=0$, with smaller values indicating outcomes that are closer to what is desired.
This abstraction recovers several important cases:
  • When $\xi=y^{\ast}$ is the global optimum of an objective function, $r(y;\xi)$ reduces to the conventional instantaneous regret in BO.
  • When no outcome-dependent pragmatic preference is imposed, as in pure BED, one may take $r(y;\xi)\equiv 0$, so that decision-making is driven entirely by epistemic value.
  • Intermediate choices recover hybrid objectives that jointly encode performance, safety, feasibility, or other task-specific desiderata.
In practice, however, the ideal regret function $r(\cdot;\xi)$ is generally unknown, implicit, or computationally intractable. A decision-maker therefore cannot act directly on $r$. Instead, it must rely on a belief-conditioned surrogate that reflects its current internal assessment of outcome desirability. This motivates the following definition of a potential energy function.
Definition 4.3 (Potential Energy Function).
The potential energy function is a belief-conditioned surrogate regret
$h_{t}:\mathcal{Y}\to\mathbb{R}_{\geq 0},\qquad h_{t}(y):=h(y\mid\mathcal{D}_{t}),$
that represents the agent’s current internal assessment of the undesirability of outcome $y$ based on the information available up to time $t$. Without loss of generality, we may shift $h_{t}$ so that $\inf_{y\in\mathcal{Y}}h_{t}(y)=0.$ Thus, $h_{t}$ serves as a time-varying, belief-dependent approximation of the true regret structure, and is updated recursively as new data are acquired.
The distinction between $r$ and $h_{t}$ is what makes PraC operational. The regret function $r$ represents the ideal task-level notion of how an outcome falls short of what is desired, whereas $h_{t}$ is the best currently available decision-usable surrogate induced by the agent’s present belief state. This mirrors the epistemic approximation used earlier in deriving PraC: just as the inaccessible true posterior is replaced by the current surrogate belief, the inaccessible true regret landscape is replaced by a belief-conditioned potential over outcomes. As more data are collected, both the predictive model and the potential energy evolve, so that the agent’s internal representation of desire is refined together with its understanding of the world.
This distinction also clarifies how $h_{t}$ should be designed in different regimes.
Known goals.
When the goal is known, $h_{t}$ can be designed to directly encode domain knowledge. In this case, the designer already has access to a meaningful notion of what constitutes shortfall, so the potential energy can be specified explicitly. Examples include classical regret with respect to an optimum, distance to a target set, penalties for constraint violation, or weighted combinations of task performance and safety requirements. In this regime, the role of $h_{t}$ is to express a known desiderative structure in a form that can be evaluated under the predictive model.
Unknown goals.
When the goal is unknown or only partially specified, the regret itself must be learned. In this regime, the pragmatic term should be endowed with additional structure, so that the agent can infer not only how the world behaves, but also what outcomes should count as desirable or undesirable. A natural construction is hierarchical: outcome desirability is mediated by additional latent variables or higher-level preference models. Given an unknown regret model $g:\mathcal{Y}\to\mathbb{R}_{\geq 0}$, one can define $h_{t}$ through an inner EFE over that regret model, yielding the nested acquisition
[…]
where $\beta_{t},\gamma_{t}\geq 0$, and $z\sim g(y)$ encodes a second-order regret, namely the regret assigned by the learned regret model itself.
In this way, the agent does not optimize with respect to a fully specified reward function given in advance; rather, it learns a surrogate of regret jointly with the predictive model. This yields a richer and more flexible account of the interplay between learning and optimization, and enables multiple layers of inference to be composed within a single decision rule. For instance, as we show later in 5.3, the model $g(y)$ can be learned through preference feedback.
This point highlights a key advantage of the PraC paradigm. Unlike many mainstream approaches that presuppose a fully specified reward function, PraC requires only a current surrogate of what can already be recognized as desirable or regrettable, together with a mechanism for refining that surrogate as evidence accumulates. One need not begin with complete illumination; it is enough to start by seeing a light in the darkness, and to let both understanding of the world and understanding of value sharpen together.
4.3 The Degree of Curiosity
If symbols specify what hidden distinctions are worth learning, and regrets specify what desired outcomes are worth pursuing, then the coefficient $\beta_{t}$ determines how much immediate pragmatic sacrifice the agent is willing to tolerate in order to refine its epistemic understanding. In this sense, $\beta_{t}$ is the exchange rate between clarifying knowledge and pursuing goals, which we denote as the degree of curiosity.
Different regimes of $\beta_{t}$ produce different modes of behavior. When $\beta_{t}=0$, the agent treats its current symbols as fixed and acts myopically with respect to the present surrogate of regret. At the other extreme, a very large $\beta_{t}$ prioritizes information acquisition even when that information has weak pragmatic relevance. Neither extreme is satisfactory: seeking knowledge without goals degenerates into aimless information gathering, whereas pursuing goals without clarified meanings risks acting on an incomplete understanding of the world.
This reveals a closed-loop structure between knowledge and action. The agent acts according to its current symbolic understanding and regret surrogate; the resulting observation then reshapes that understanding and changes which future actions become desirable. Thus, curiosity should not be viewed as a static exploration heuristic. Rather, it is a regulated gain in this knowledge–action loop: it determines how strongly unresolved uncertainty should influence action when the agent’s current understanding may still be insufficient for reliable decision-making. This motivates designing $\beta_{t}$ as an adaptive controller for the knowledge–action loop.
Feedforward scale.
The feedforward component estimates the scale of curiosity required for epistemic value to be comparable with pragmatic value. Given a finite candidate set $\mathcal{X}_{t}^{\mathrm{cand}}$, let
$\widehat{I}_{t}(x)\approx I(s;(x,y)\mid\mathcal{D}_{t}),\qquad\widehat{H}_{t}(x):=\mathbb{E}[h_{t}(y)\mid x,\mathcal{D}_{t}],$
where $\widehat{I}_{t}(x)$ estimates the epistemic value of querying $x$, and $\widehat{H}_{t}(x)$ estimates its expected pragmatic regret. To avoid unstable ratios when information gain is negligible, define the informative candidate set
$\mathcal{X}_{t}^{\mathrm{info}}=\{x\in\mathcal{X}_{t}^{\mathrm{cand}}:\widehat{I}_{t}(x)>\varepsilon_{I}\},$
where $\varepsilon_{I}>0$ is a numerical floor. If $\mathcal{X}_{t}^{\mathrm{info}}$ is nonempty, we define the feedforward curiosity scale as
$\beta_{t}^{\mathrm{ff}}=\operatorname{Quantile}_{x\in\mathcal{X}_{t}^{\mathrm{info}}}\frac{\widehat{H}_{t}(x)}{\widehat{I}_{t}(x)+\varepsilon_{I}}.$
If $\mathcal{X}_{t}^{\mathrm{info}}=\varnothing$, we set $\beta_{t}^{\mathrm{ff}}=\beta_{\min}$, since the current candidate set contains no sufficiently informative query.
This feedforward term $\beta_{t}^{\mathrm{ff}}$ estimates the local exchange rate between information and regret: when useful information is scarce relative to pragmatic regret, the required curiosity scale increases; when information is readily available, a smaller scale is sufficient to make information compete with immediate pragmatic value.
Feedback activation.
The feedback component determines whether epistemic pressure is still needed. Let $U_{t}$ be a scalar uncertainty measure over the relevant symbol, such as posterior entropy, and let $U^{\star}$ denote the desired epistemic equilibrium, typically zero or an irreducible uncertainty floor. Define
$\beta_{t}^{\mathrm{fb}}=k_{\beta}\operatorname{clip}\left(\frac{U_{t}-U^{\star}}{U_{0}-U^{\star}+\varepsilon_{U}},0,1\right),$
where $U_{0}$ is the initial uncertainty, $k_{\beta}$ is a feedback gain, and $\varepsilon_{U}>0$ prevents numerical instability.
The combined schedule for curiosity coefficient is then
$\beta_{t}=\operatorname{clip}\left(\beta_{t}^{\mathrm{fb}}\beta_{t}^{\mathrm{ff}},\beta_{\min},\beta_{\max}\right),$
where the two components play complementary roles: the feedforward term estimates how large curiosity must be for information to compete with pragmatic value, while the feedback activation suppresses curiosity as the belief approaches the desired epistemic equilibrium. Thus, curiosity is high only when information is both decision-relevant and epistemically needed; it relaxes when the relevant uncertainty has been resolved.